Solve using elimination method.

2x+ 8y = 26
5x- 7y = -16

Answer by faceoff57(108)

(2x+ 8y = 26)-5
(5x- 7y = -16)2
-10x-40y=-130
10x-14y=- 32
-54y=-162
y= 3
2x+8(3)=26
2x+24=26
2x=2
x=1